OpenStudy (anonymous):

what are the solutions of the inequality? x – 3 ≤ –12

OpenStudy (anonymous):

\[x-3{\le}-12\]\[x-3+3{\le}-12+3\]\[x{\le}-9\]Thus the answer is \[[-9, \infty)\]

OpenStudy (anonymous):

sorry it is \[(-{\infty},-9]\]

OpenStudy (anonymous):

okay thank you! what about q – 1/2>1/3

OpenStudy (anonymous):

\[q-\frac{ 1 }{ 2 }>\frac{ 1 }{ 3 }\]\[q>\frac{ 1 }{ 3 }+\frac{ 1 }{ 2 }\]\[q>\frac{ 5 }{ 6 }\]